# Energy based diffusion generator for efficient sampling of Boltzmann distributions

### v2

@@ Abstract
Sampling from Boltzmann distributions, particularly those tied to high-dimensional and complex energy functions, poses a significant challenge in many fields. In this work, we present the Energy-Based Diffusion Generator (EDG), a novel approach that integrates ideas from variational autoencoders and diffusion models. EDG leverages a decoder to transform latent variables from a simple distribution into samples approximating the target Boltzmann distribution, while the diffusion-based encoder provides an accurate estimate of the Kullback-Leibler divergence during training. Notably, EDG is simulation-free, eliminating the need to solve ordinary or stochastic differential equations during training. Furthermore, by removing constraints such as bijectivity in the decoder, EDG allows for flexible network design. Through empirical evaluation, we demonstrate the superior performance of EDG across a variety of complex distribution tasks, outperforming existing methods.
@@ 1 Introduction
In various fields such as computational chemistry, statistical physics, and machine learning, the challenge of sampling from a Boltzmann distribution corresponding to a high-dimensional and complex energy function is ubiquitous [1, 2]. Unlike training tasks for data-driven generative models, where pre-sampled data can be utilized to learn complex distributions, sampling from Boltzmann distributions presents a unique and significant challenge due to the lack of readily available data [3, 4]. For example, simulating the phase transition of the Ising model can be framed as a sampling problem given the energy function, which presents a complex and difficult problem that has yet not to be effectively addressed [5, 6].
Markov Chain Monte Carlo (MCMC) methods [7], along with Brownian and Hamiltonian dynamics [8, 9, 10, 11], have offered a pivotal solution to the challenge of sampling from high-dimensional distributions. These methods operate by iteratively generating candidates and updating samples, ultimately achieving asymptotic unbiasedness at the limit of infinite sampling steps. In recent years, researchers have proposed adaptive MCMC as a strategy for generating candidate samples, showcasing notable advancements in augmenting the efficiency and effectiveness of the sampling process [12, 13, 14]. However, the prolonged mixing time of MCMC still constrains its performance. Several studies have shown that using neural networks to construct and optimize proposal distributions in MCMC can significantly improve its efficiency [13, 15, 16]. However, there is still a lack of an effective and widely adaptable loss function to facilitate this optimization.
Variational inference (VI) is another crucial approach for addressing intractable distributions. VI leverages a generator capable of swiftly producing samples to approximate the target Boltzmann distribution, and the optimization of the generator’s parameters is then undertaken to minimize a statistical distance, such as the Kullback-Leibler (KL) divergence, between the distribution of the generated samples and the target distribution. Due to its ability to model complex distributions and provide explicit probability density functions, normalizing flow (NF) has been extensively applied to construct generators for VI methods [17, 18, 19, 20, 21, 22, 23, 24, 25, 26]. However, the bijective nature of NFs imposes a constraint on their effective capacity, often rendering it insufficient for certain sampling tasks. Given the target density function and the generated samples, Stein Discrepancy [27, 28] provides a different means to evaluate the goodness of fit, while the computation of kernel functions and their gradients limits its performance in high-dimensional tasks. In addition, the combination of MCMC and VI methods stands as a current focal point in research [29, 30, 31, 32, 33, 34]. This combination seeks to harness the strengths of both approaches, offering a promising avenue for addressing the challenges associated with sampling from high-dimensional distributions and enhancing the efficiency of probabilistic modeling.
@@
With the prosperity of diffusion based generative models [35, 36, 37, 38], they have been applied to address challenges in the sampling problem. By training time-dependent score matching neural networks, methods proposed in [39, 40, 41] shape the Gaussian distribution into complex target densities, employing the KL divergence as the loss function. To mitigate mode-seeking issues, [42] introduces the log-variance loss, showcasing favorable properties. Additionally, an alternative training objective is outlined in [43], relying on flexible interpolations of the energy function and demonstrating substantial improvements for multi-modal targets. However, a common drawback of these methods is their reliance on numerical differential equation solvers for computing time integrals, which can lead to substantial computational costs.
@@
In this research endeavor, we present a novel approach termed the energy based diffusion generator (EDG), drawing inspiration from both the variational autoencoder (VAE) technique [44] and the diffusion model. The architecture of EDG closely resembles that of VAE, comprising a decoder and an encoder. The decoder is flexible in mapping latent variables distributed according to tractable distributions to samples, without the imposition of constraints such as bijectivity, and we design in this work a decoder based on generalized Hamiltonian dynamics to enhance sampling efficiency. The encoder utilizes a diffusion process, enabling the application of score matching techniques for precise and efficient modeling of the conditional distribution of latent variables given samples. Unlike existing diffusion-based methods, the loss function of EDG facilitates the convenient computation of unbiased estimates in a random mini-batch manner, removing the need for numerical solutions to ordinary differential equations (ODEs) or stochastic differential equations (SDEs) during training. Numerical experiments conclusively demonstrate the effectiveness of EDG.
2 Preliminaries and setup
@@
where the normalizing constant $Z=\int\exp(-U(x))\mathrm{d}x$ is usually computationally intractable. To tackle this challenge, the Boltzmann generator [18], along with its various extensions [24, 25, 26], has emerged as a prominent technique in recent years. These approaches harness NF to parameterize a trainable and analytical density function, with parameter optimization achieved through the minimization of the KL divergence between the surrogate density and $\pi$. Nevertheless, in contrast to typical generative models, the pursuit of exact probability density computation imposes substantial constraints on NFs: Each transformation layer must be a bijection, and the determinant of its Jacobian matrix can be tractably computed. These requirements inherently limit the capacity of NFs to effectively model complex distributions.
@@
It is important to note that, unlike conventional data-driven VAEs, we do not have access to samples from the target distribution $\pi(x)$. In fact, obtaining such samples is precisely the goal of the generator. As a result, the variational approximation of the KL divergence $D_{KL}\left(\pi(x)||p_{D}(x)\right)$ cannot be used to train the model. Instead, in this work, we consider the following divergence and its upper bound:
$D_{KL}\left(p_{D}(x)||\pi(x)\right)$$\leq$ $D_{KL}\left(p_{D}(z_{0})\cdot p_{D}(x|z_{0},\phi)||\pi(x)\cdot p_{E}(z_{0}|x,\theta)\right)$ (1)
$=$ $\mathbb{E}_{p_{D}(z_{0})\cdot p_{D}(x|z_{0},\phi)}\left[\log\frac{p_{D}(z_{0})p_{D}(x|z_{0},\phi)}{p_{E}(z_{0}|x,\theta)}+U(x)\right]$
$+\log Z.$
Here, the parametric distribution $p_{E}(z_{0}|x,\theta)$ defines an encoder that maps from $x$ to the latent variable $z$, and the equality is achieved if $p_{E}(z|x,\theta)$ matches the conditional distribution of $z$ for a given $x$, deduced from the decoder.
It seems that we have only increased the complexity of the problem, as we are still required to approximate a conditional distribution. However, in the upcoming section, we demonstrate that we can effectively construct the encoder using the diffusion model [35, 37] and optimize all parameters without the need to numerically solve ordinary or stochastic differential equations.
@@ 3 Energy-based diffusion generator
The diffusion model [37, 38] has emerged in recent years as a highly effective approach for estimating data distributions. Its core idea is to construct a diffusion process that progressively transforms data into simple white noise, and learn the reverse process to recover the data distribution from noise. In this work, we apply the principles of the diffusion model by incorporating a diffusion process into the latent space, enabling us to efficiently overcome the challenges arising from the variational framework for the sampling problem defined by Equation (1). We refer to the model produced by this method as the energy-based diffusion generator (EDG).
@@ 3.1 Model architecture
In the EDG framework, we initiate a diffusion process from the latent variable $z_{0}$ and combine it with the decoder, resulting in what we term the “decoding process”:
$z_{0}\in R^{D}\sim p_{D}(z_{0})\triangleq\mathcal{N}(x|0,I),\quad x|z_{0}\sim p_{D}(x|z_{0};\phi)$
$\mathrm{d}z_{t}=f(z_{t},t)\mathrm{d}t+g(t)\mathrm{d}W_{t},\quad t\in[0,T]$ (2)
where $W_{t}$ is the standard Wiener process, $f(\cdot,t):R^{D}\rightarrow R^{D}$ acts as the drift coefficient, and $g(\cdot):R\rightarrow R$ serves as the diffusion coefficient. To streamline notation, we denote the probability distribution defined by the decoding process as $p_{D}$. In typical SDEs applied in diffusion models, two critical conditions hold: (a) the transition density $p_{D}(z_{t}|z_{0})$ can be analytically computed without numerically solving the Fokker-Planck equation, and (b) $z_{T}$ is approximately uninformative with $p_{D}(z_{T})\approx p_{D}(z_{T}|z_{0})$.
If we only consider the statistical properties of the latent diffusion process $z_{[\cdot]}=\{z_{t}\}_{t\in[0,T]}$, it is uninformative and only describes a transition from one simple noise to another. However, when we account for the conditional distribution of $z_{t}$ given a sample $x$, the process $z_{[:]}$ represents the gradual transformation of the complex conditional distribution $p_{D}(z_{0}|x)\propto p_{D}(z_{0})\cdot p_{D}(x|z_{0})$ into the tractable distribution $p_{D}(z_{T}|x)=p_{D}(z_{T})$, where the independence between $z_{T}$ and $x$ results from the independence between $z_{0}$ and $z_{T}$ (see A). This implies that, starting from $z_{T}\sim p_{D}(z_{T})$, we can obtain samples from $p_{D}(z_{0}|x)$ by simulating the following reverse-time diffusion equation [45]:
$\mathrm{d}z_{\tilde{t}}=-\left(f(z_{\tilde{t}},\tilde{t})-g(\tilde{t})^{2}\nabla_{z_{\tilde{t}}}\log p_{D}(z_{\tilde{t}}|x)\right)\mathrm{d}\tilde{t}+g(\tilde{t})\mathrm{d}W_{\tilde{t}},$ (3)
where $\tilde{t}=T-t$ denotes the reverse time. As in conventional diffusion models, practical implementation of this simulation is challenging due to the intractability of the score function $\nabla_{z_{\tilde{t}}}\log p_{D}(z_{\tilde{t}}|x)$, and we therefore also use a neural network to approximate the score function, denoted as $s(z_{\tilde{t}},x,\tilde{t};\theta)$. This approximation leads to what we refer to as the “encoding process”, achieved by integrating the parametric reverse-time diffusion process and the target distribution of $x$:
@@
$\mathrm{d}z_{\tilde{t}}=-\left(f(z_{\tilde{t}},\tilde{t})-g(\tilde{t})^{2}s(z_{\tilde{t}},x,\tilde{t};\theta)\right)\mathrm{d}\tilde{t}+g(\tilde{t})\mathrm{d}W_{\tilde{t}},\quad\tilde{t}=T-t.$ (4)

### v3

# Energy-based diffusion generator for efficient sampling of Boltzmann distributions

@@ Abstract
Sampling from Boltzmann distributions, particularly those tied to high dimensional and complex energy functions, poses a significant challenge in many fields. In this work, we present the Energy-Based Diffusion Generator (EDG), a novel approach that integrates ideas from variational autoencoders and diffusion models. EDG uses a decoder to generate Boltzmann-distributed samples from simple latent variables, and a diffusion-based encoder to estimate the Kullback-Leibler divergence to the target distribution. Notably, EDG is simulation-free, eliminating the need to solve ordinary or stochastic differential equations during training. Furthermore, by removing constraints such as bijectivity in the decoder, EDG allows for flexible network design. Through empirical evaluation, we demonstrate the superior performance of EDG across a variety of sampling tasks with complex target distributions, outperforming existing methods.
@@ 1 Introduction
In various fields such as computational chemistry, statistical physics, and machine learning, the challenge of sampling from a Boltzmann distribution corresponding to a high-dimensional and complex energy function is ubiquitous [rapaport2004art]. Unlike training tasks for data-driven generative models, where pre-sampled data can be utilized to learn complex distributions, sampling from Boltzmann distributions presents a unique and significant challenge due to the lack of readily available data [martino2013flexibility, rubinstein2016simulation]. For example, simulating the phase transition of the Ising model can be framed as a sampling problem given the energy function, which presents a complex and difficult problem that has yet not to be effectively addressed [4, 5].
Markov Chain Monte Carlo (MCMC) methods [mc-1], along with Brownian and Hamiltonian dynamics [mc-2, araki2013adaptive, araki2015efficient, mc-6], have offered a pivotal solution to the challenge of sampling from high-dimensional distributions. These methods operate by iteratively generating candidates and updating samples, ultimately achieving asymptotic unbiasedness at the limit of infinite sampling steps. Meanwhile, some enhanced sampling methods [vanden2012rare, martinsson2019simulated] developed for rare event problems accelerate sampling efficiency while significantly reducing relative statistical errors, even for low-probability regions.
In recent years, researchers have proposed adaptive MCMC as a strategy for generating candidate samples, showcasing notable advancements in augmenting the efficiency and effectiveness of the sampling process [mc-4, mc-5, galliano2024policy]. However, designing optimal loss functions is always challenging for the sampling problems. Many distributional discrepancies used in generative models, such as Jensen-Shannon divergence, maximum mean discrepancy and Wasserstein distance, require access to samples from the true distribution, which are unavailable in our setting. Stein discrepancy derived from Stein’s identity [vi-8], and the Kullback-Leibler (KL) divergence, between the distribution of the generated samples and the target distribution, may serve as potential alternatives for sampling.
Variational inference (VI) is a widely used approach to sampling problems. It employs a generator that efficiently produces samples to approximate the target Boltzmann distribution, and optimizes the generator’s parameters using the KL divergence, without requiring true samples from the target distribution. Due to its ability to model complex distributions and provide explicit probability density functions, normalizing flow (NF) [dinh2014nice, dinh2016density, chao2023training, nijkamp2020mcmc] has been extensively applied to construct generators for VI methods [key-4, vi-2]. Furthermore, a growing body of research has expanded their applicability across diverse domains [dibak2022temperature, kohler2021smooth, ising, wirnsberger2023estimating, schebek2024efficient, coretti2025learning]. However, the bijective nature of NFs imposes a constraint on their effective capacity, often rendering it insufficient for certain sampling tasks. Continuous normalizing flows [klein2024transferable, kohler2020equivariant, albergo2022building, jung2024normalizing] represent a specific subclass of NF in which the invertible transformation is defined through an ordinary differential equation. For more relevant studies, please refer to the Related Work in Sec. 2.1.
@@
With the prosperity of diffusion based generative models [diff-2, key-2, key-3, bou2014metropolis], they have been applied to address challenges in the sampling problem. By training time-dependent score matching neural networks, methods proposed in [diff-6, diff-4, diff-3] shape the Gaussian distribution into complex target densities, employing the KL divergence as the loss function. [diff-1] also explores multiple connections between diffusion and other VI methods. To mitigate mode-seeking issues, [diff-7] introduces the log-variance loss, showcasing favorable properties. Additionally, an alternative training objective is outlined in [diff-8], relying on flexible interpolations of the energy function and demonstrating substantial improvements for multi-modal targets. However, a common drawback of these methods is their reliance on numerical differential equation solvers for computing time integrals, which can lead to substantial computational costs.
@@
In this research endeavor, we present a novel approach termed the energy-based diffusion generator (EDG), drawing inspiration from both the variational autoencoder (VAE) technique [kingma2019introduction] and the diffusion model. The architecture of EDG closely resembles that of VAE, comprising a decoder and an encoder. The decoder is flexible in mapping latent variables distributed according to tractable distributions to samples, without the imposition of constraints such as bijectivity, and we design in this work a decoder based on generalized Hamiltonian dynamics to enhance sampling efficiency. The encoder utilizes a diffusion process, enabling the application of score matching techniques for precise and efficient modeling of the conditional distribution of latent variables given samples. Unlike existing diffusion-based sampling methods, the loss function of EDG facilitates the convenient computation of unbiased estimates in a random mini-batch manner, removing the need for numerical solutions to ordinary differential equations (ODEs) or stochastic differential equations (SDEs) during training. Numerical experiments conclusively demonstrate the effectiveness of EDG.
2 Related work and preliminaries
2.1 Related work
Variational inference-based sampling algorithms: The Boltzmann generator (BG) [bg3, vi-2] is one of the most representative sampling methods based on VI. It leverages NFs to parameterize a trainable and analytically tractable density, with parameter optimization carried out by minimizing the KL divergence between the surrogate and target distributions. Here, NFs are composed of stacked bijective transformations, allowing the density to be computed in closed form. In addition, the combination of MCMC and VI methods stands as a current focal point in research [vi-3, vi-4, vi-6, vi-5, shen2021randomizing]. This combination seeks to harness the strengths of both approaches, offering a promising avenue for addressing the challenges associated with sampling from high-dimensional distributions and enhancing the efficiency of probabilistic modeling.
Adaptive MCMC: Recent work has augmented MCMC sampling with nonlocal transition kernels parameterized by neural networks, including the NF-based method proposed by Vanden-Eijnden et al. [gabrie2022adaptive] and the diffusion model-based method [akhound2024iterated]. This synergistic approach enables neural networks to accelerate MCMC sampling while the model are trained in a data-driven mode from samples of MCMC.
Stein-based models: Designing optimal loss functions for sampling problems is always challenging for the sampling problems. Stein discrepancy has emerged as a promising candidate [vi-7, vi-8, grathwohl2020learning, di2021neural], as it allows direct evaluation of the discrepancy between generated samples and the target distribution, without requiring real samples from the target or access to the density of the generated distribution. In particular, the kernelized Stein discrepancy [vi-8] offers a practical and convenient criterion, which can test the goodness of fit easily. The trainable Stein Discrepancy [grathwohl2020learning, di2021neural] further leverages neural networks to parameterize a class of functions based on the Stein’s Identity, using the learned discrepancy to enhance its capability.
Latent diffusion models: Latent diffusion models, which are primarily designed to address generative modeling in data-driven scenarios, have recently gained significant attention [rombach2022high, fu2024latent, zheng2024ld]. Their main idea to use a pre-trained encoder and decoder to obtain a latent space that both effectively represents the data and facilitates efficient sampling, with the diffusion model learning the distribution of latent variables. However, its integration with sampling problems remains to be explored.
2.2 Preliminaries and Setup
@@
where the normalizing constant $Z=\int\exp(-U(x))\mathrm{d}x$ is usually computationally intractable. To tackle this challenge, the BG [key-4], along with its various extensions [bg2, bg3, bg4], has emerged as a prominent technique in recent years. In this work, we aim to overcome the limitations of BG by employing a generator with fewer structural constraints (e.g., without requiring bijection), and design a flexible model tailored for sampling tasks.
@@
It is important to note that, unlike conventional data-driven VAEs, we do not have access to samples from the target distribution $\pi(x)$. In fact, obtaining such samples is precisely the goal of the generator. As a result, the variational approximation of the KL divergence $D_{\mathrm{KL}}\left(\pi(x)||p_{D}(x)\right)$ cannot be used to train the model. Instead, in this work, we consider the divergence $D_{\mathrm{KL}}\left(p_{D}(x)||\pi(x)\right)$ and its upper bound:
$D_{\mathrm{KL}}\left(p_{D}(x)||\pi(x)\right)$ $\leq D_{\mathrm{KL}}\left(p_{D}(z_{0})p_{D}(x|z_{0};\phi)||\pi(x)p_{E}(z_{0}|x;\theta)\right)$
$=\mathbb{E}_{p_{D}(z_{0})p_{D}(x|z_{0};\phi)}\left[\log\frac{p_{D}(z_{0})p_{D}(x|z_{0};\phi)}{p_{E}(z_{0}|x;\theta)}+U(x)\right]$
$\qquad+\log Z.$ (1)
Here, the parametric distribution $p_{E}(z_{0}|x;\theta)$ defines an encoder that maps from $x$ to the latent variable $z_{0}$, and the equality is achieved if $p_{E}(z_{0}|x;\theta)$ matches the conditional distribution of $z$ for a given $x$, deduced from the decoder. The detailed proof is provided in A.
It seems that we have only increased the complexity of the problem, as we are still required to approximate a conditional distribution. However, in the upcoming section, we demonstrate that we can effectively construct the encoder using the diffusion model [diff-1, key-2] and optimize all parameters without the need to numerically solve ordinary or stochastic differential equations.
@@ 3 Energy-based diffusion generator
The diffusion model [key-2, key-3] has emerged in recent years as a highly effective approach for estimating data distributions. Its core idea is to construct a diffusion process that progressively transforms data into simple white noise, and learn the reverse process to recover the data distribution from noise. In this work, we apply the principles of the diffusion model by incorporating a diffusion process into the latent space, enabling us to efficiently overcome the challenges arising from the variational framework for the sampling problem defined by Eq. (1). We refer to the model produced by this method as the energy-based diffusion generator (EDG).
@@ 3.1 Model architecture
In the EDG framework, we initiate a diffusion process from the latent variable $z_{0}$ and combine it with the decoder, resulting in what we term the “decoding process” $p_{D}$
$z_{0}\in\mathbb{R}^{d}\sim p_{D}(z_{0})\triangleq\mathcal{N}(x|0,I),\quad x|z_{0}\sim p_{D}(x|z_{0};\phi)$
$\mathrm{d}z_{t}=f(z_{t},t)\mathrm{d}t+g(t)\mathrm{d}{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}W_{t}},\quad t\in[0,T]$ (2)
where $W_{t}$ is the standard Wiener process, $f(\cdot,t):\mathbb{R}^{d}\rightarrow\mathbb{R}^{d}$ acts as the drift coefficient, and $g(\cdot):R\rightarrow R$ serves as the diffusion coefficient. To streamline notation, we denote the probability distribution defined by the decoding process as $p_{D}$, and represent the variables defined by the SDE as $z_{[\cdot]}=\{z_{t}\}_{t\in[0,T]}$. In typical SDEs applied in diffusion models, two critical conditions hold: (a) the transition density $p_{D}(z_{t}|z_{0})$ can be analytically computed without numerically solving the Fokker-Planck equation, and (b) $z_{T}$ is approximately uninformative with $p_{D}(z_{T})\approx p_{D}(z_{T}|z_{0})$.
If we only consider the statistical properties of the latent diffusion process, it is uninformative and only describes a transition from one simple noise to another. However, when we account for the conditional distribution of $z_{t}$ given a sample $x$, the process $z_{[\cdot]}$ represents the gradual transformation of the complex conditional distribution $p_{D}(z_{0}|x)\propto p_{D}(z_{0})\cdot p_{D}(x|z_{0})$ into the tractable distribution $p_{D}(z_{T}|x)=p_{D}(z_{T})$, where the independence between $z_{T}$ and $x$ results from the independence between $z_{0}$ and $z_{T}$ (see B). This implies that, starting from $z_{T}\sim p_{D}(z_{T})$, we can obtain samples from $p_{D}(z_{0}|x)$ by simulating the following reverse-time diffusion equation [diff-9]:
$\mathrm{d}z_{\tilde{t}}=-\left(f(z_{\tilde{t}},\tilde{t})-g(\tilde{t})^{2}\nabla_{z_{\tilde{t}}}\log p_{D}(z_{\tilde{t}}|x)\right)\mathrm{d}\tilde{t}+g(\tilde{t})\mathrm{d}{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\widetilde{W}_{\tilde{t}}}$, (3)
where $\tilde{t}=T-t$ denotes the reverse time. The process $\widetilde{W}$ is a distinct standard Wiener process, whose non-trivial relationship with $W$ is formally discussed in [diff-9]. As in conventional diffusion models, practical implementation of this simulation is challenging due to the intractability of the score function $\nabla_{z_{\tilde{t}}}\log p_{D}(z_{\tilde{t}}|x)$, and we therefore also use a neural network to approximate the score function, denoted as $s(z_{\tilde{t}},x,\tilde{t};\theta)$. This approximation leads to what we refer to as the “encoding process”, achieved by integrating the parametric reverse-time diffusion process and the target distribution of $x$:
@@
$\mathrm{d}z_{\tilde{t}}=-\left(f(z_{\tilde{t}},\tilde{t})-g(\tilde{t})^{2}s(z_{\tilde{t}},x,\tilde{t};\theta)\right)\mathrm{d}\tilde{t}+g(\tilde{t})\mathrm{d}{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\widetilde{W}_{\tilde{t}}},\quad\tilde{t}=T-t.$ (4)